OpenStudy (anonymous):

find the equation of the tangent line at the give point on the curve. xy + x = 2 (1,1)

OpenStudy (anonymous):

take the derivative get \[y+xy'+1=0\] replace \(x\) and \(y\) by 1, solve for \(y'\) to get the slope then use the point slope formula

OpenStudy (anonymous):

alright thank you.

OpenStudy (anonymous):

so the derivative would be : y'=(-y-1)/x y'(1,1)=-1-1/1 y'(1,1)=-2 so slope would be : y-y1=mtan(x-x1) y-1=-2(x-1) y=-2x+3 Is this correct?